OpenStudy (anonymous):

Given the law of sines, all the measures of a triangle can be found with two sides and any angle. True/False

OpenStudy (anonymous):

False.

OpenStudy (precal):

well you must know one angle and side that goes together \[\frac{a}{sinA}=\frac{b}{sinB}=\frac{c}{sinC}\]